OpenStudy (anonymous):

Eileen saves dimes and quarters. She has 40 coins which totals $6.55, how many of each coins does she has? (has to be a linear equation)

OpenStudy (anonymous):

x+y = 40

OpenStudy (texaschic101):

d + q = 40 -->d = 40 - q 0.10d + 0.25q = 6.55 now sub 40 - q in for d in the 2nd equation.. 0.10(40 - q) + 0.25q = 6.55 4.00 - 0.10q + 0.25q = 6.55 4.00 + 0.15q = 6.55 0.15q = 6.55 - 4.00 0.15q = 2.55 q = 2.55/0.15 q = 17 now sub 17 in for q in the 1st equation... d + q = 40 d + 17 = 40 d = 40 - 17 d = 23 check... 0.10d + 0.25q = 6.55 0.10(23) + 0.25(17) = 6.55 2.30 + 4.25 = 6.55 6.55 = 6.55 (correct) ANSWER : Ellen has 17 quarters totaling 4.25 and 23 dimes totaling 2.30. Together there are 40 quarters and dimes totaling 6.55.